Ticket Number: HEA-NOC /20020516-01 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: EPA
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start:
Ticket Closed: 20020524 09:30:21 UTC Problem End:
EPA are upgrading their Firewall on Friday the 16th of May.
All connectivity to the EPA.
Mail will be backed up by heanet
EPA report that the upgrade has been completed
Time to Fix:
Outage expected for all of tomorrow, starting in mid morning